Circular No. 7736

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ez IN PGC 17642 D. J. Hutchings and W. D. Li, University of California at Berkeley

(UCB), report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope on Oct. 17.5 (mag ∼ 17.8) and 18.5 UT (mag ∼ 17.6). SN 2001ez is located at α = 5^{h}42^{m}37^{s}.64, δ = +69^{o}14′14′′.6 (equinox 2000.0), which is 3′′.3 east and 2′′.0 north of the nucleus of PGC 17642. A KAIT image taken on Oct. 2.5 shows nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2001ev IN UGC 2653 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ev (cf. IAUC 7732), obtained by M. Calkins on Oct. 17.39 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 6939 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hα line is 11 800 km/s.

SUPERNOVA 2001ex IN UGC 3595 Matheson et al. also report that a spectrum of SN 2001ex (cf. IAUC

7735), obtained by M. Calkins as above on Oct. 18.50 UT, shows it to be a type-Ia supernova near maximum light. Adopting the NED recession velocity of 7908 km/s for the host galaxy, the supernova expansion velocity is ∼ 9800 km/s for Si II (rest 635.5 nm).

SUPERNOVA 2001ey IN MCG –01-57-10 Matheson et al. add that a spectrum (range 370–750 nm) of SN 2001ey

(cf. IAUC 7735), obtained by M. Calkins as above on Oct. 18.08 UT, shows it to be a type-IIn supernova. Strong hydrogen Balmer emission lines are superposed on a nearly featureless continuum.

W. M. Wood-Vasey, UCB and Lawrence Berkeley National Laboratory, reports that observations at the location of 2001ey taken with the Oschin 1.22-m Schmidt telescope at Palomar on Oct. 5.13 UT show nothing to a limiting unfiltered magnitude of 19.2.

2001 October 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green